OpenStudy (anonymous):

Solve: 9x^2 - 16x + 60 = 0 On my homework it says solve the equation by determining the discriminant. And then show your work. If you can tell me how to do it and I can find the answer, that would be awesome.

OpenStudy (anonymous):

9x^2-16+60=0

OpenStudy (anonymous):

(-16+60)=44

OpenStudy (anonymous):

9x^2+44=0

OpenStudy (anonymous):

-44 from both sides

OpenStudy (anonymous):

9x^2=-44

OpenStudy (anonymous):

divide both sides by 9

OpenStudy (anonymous):

Dang You wrote fast.

OpenStudy (anonymous):

x^2=-44/9

OpenStudy (anonymous):

sqr both sides

OpenStudy (anonymous):

x=sqr(-44/9)

OpenStudy (anonymous):

x= I(sqr(44/9))

OpenStudy (anonymous):

does that sound about right?

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

What is with the 44?

OpenStudy (anonymous):

-16+60=44

OpenStudy (anonymous):

**** oh no its a -16x

OpenStudy (anonymous):

lol so freaking stupid xD

OpenStudy (anonymous):

oh then that changes everything

OpenStudy (anonymous):

9x^2-16x+60=0

OpenStudy (anonymous):

Yeah so sorry

OpenStudy (anonymous):

I cant find a rational solution sorry but try this http://www.wolframalpha.com/input/?i=9x%5E2-16x%2B60%3D0

OpenStudy (anonymous):

Lol im on the same thing right now looking at it thanks tho! :)

OpenStudy (anonymous):

sorry :)